OpenStudy (anonymous):

Reduce the following rational expression to lowest terms, if possible. (Problem in comments)

OpenStudy (anonymous):

\[\frac{ 7x+21 }{ 7x^2+42x+63 }\]

hartnn (hartnn):

first factor 7 from the denominator and numerator then, it will be easier to factor the quadratic in denominator :)

OpenStudy (anonymous):

\[\frac{ 7(x+3) }{ 7(x+3)^2 }\] two of the 7(x+3)s cancel out & im left with one on the bottom sooooo... \[\frac{ 1 }{ 7(x+3) }\]

OpenStudy (anonymous):

Internet went out...

hartnn (hartnn):

the 7 also cancels out from numerator and denominator!

OpenStudy (anonymous):

So only the x+3 is a square? not the 7? oh ok

hartnn (hartnn):

\(\huge \dfrac{ \cancel 7\cancel {(x+3)} }{\cancel 7(x+3)^\cancel 2 }\)

OpenStudy (anonymous):

Wait, duh...what am i saying lol i got it now